two boats A and B move away from a buoy anchored at the middle of a river along mutually perpendicular straight lines. The boat A moves along the stream and the boat B across the river. After moving off an equal distance from the buot, both the boats returned to their original position. Calculate the ratio of the times taken by boat A to that taken by boat B if the velocity of each boat with respect to still water is 2 times the stream velocity

Grade:11

1 Answers

Aman Bansal
12 years ago

Dear Aditi,

et u be the stream velocity, then velocity of each boat(with respect to water) is v = 1.2 u.

Moving away from buoy : If A moves away a distance d in time {t_A} and B in {t_B}, then

{T_B} = \frac{d}{{{v_B}}} = \frac{d}{{v\cos \theta }} = \frac{d}{{v.\frac{{\sqrt {{v^2} - {u^2}} }}{v}}}

 

\;\,\, = \frac{d}{{\sqrt {{v^2} - {u^2}} }}\left[ {{\bf{since}}{\rm{ }}\;v\sin \theta  = u\,\;\;\;{\rm{i}}{\rm{.e}}{\rm{.}}\;\;\sin \theta  = u{\rm{/}}v} \right]
and\;\;\;\;\;{t_A} = \frac{d}{{v + u}}

Coming back to buoy: If A takes a time of {t_A}' seconds and B takes {t_B}' seconds in coming back to buoy, then+

t_B^' = \frac{d}{{\sqrt {{v^2} - {u^2}} }}
and\;\;\;\;\;t_A^' = \frac{d}{{v - u}}

Therefore, total time of motion of B is

Hence\;\;\;\;\;\frac{{{T_A}}}{{{T_B}}} = \frac{{1.2}}{{\sqrt {{{(1.2)}^2} - 1} }} = 1.8

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

 

{T_B} = {t_B} + {t_B}' = 2{t_B} = \frac{{2d}}{{\sqrt {{v^2} - {u^2}} }}

and that for A is

{T_A} = {t_A} + {t_A}' = \frac{d}{{v + u}} + \frac{d}{{v - u}}
\;\,\, = \frac{{2vd}}{{{v^2} - {u^2}}}